OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (kidrah69):

Thats wrong...

OpenStudy (anonymous):

now im confused

OpenStudy (kidrah69):

ok lets use sustitution.

OpenStudy (kidrah69):

we will sub y=2x+1 into equation 2

OpenStudy (kidrah69):

-4x+2x+1=-1 combine like terms

OpenStudy (anonymous):

2x+1=-1

OpenStudy (kidrah69):

careful it would be -2x+1=-1 now isolate x

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

x+1=-2

OpenStudy (anonymous):

x=-1?

OpenStudy (kidrah69):

yep :)

OpenStudy (anonymous):

whats next?

OpenStudy (kidrah69):

now we plug -1 into our equation 1 y=2(-1)+1

OpenStudy (anonymous):

y=-1?

OpenStudy (kidrah69):

we made a mistake somewhere let me double check

OpenStudy (kidrah69):

Equation 1 ) y=2x+1 Equation 2) -4x+y=-1 -4x+(2x+1)=-1 -1 is right here :) y=-2-1 y=-3

OpenStudy (anonymous):

oh ok thanks for the help man!

OpenStudy (kidrah69):

No probs :)

OpenStudy (kidrah69):

so it would be (1,3)

OpenStudy (kidrah69):

the negative 1 on top should be postive it was a typo sorry :)